# HunyuanWorld 1.0: Generating Immersive, Explorable, and Interactive 3D Worlds from Words or Pixels

### v1

@@ 2.1 Generating Panoramas As World Proxy
Panorama Generation. The architecture of our panorama generation model (Panorama-DiT) is based on the Diffusion Transformer (DiT) framework [35]. For text-to-panorama generation, only the enhanced text prompts are fed into the diffusion model as condition. For image-to-panorama generation, we first project the input image into panoramic space and then encode it into the latent space using a variational autoencoder (VAE). Next, we concatenate the condition image with the noisy latent for diffusion model. To enhance the generation quality and provide additional control, the image-to-panorama generation process is also conditioned on an auxiliary textual description produced by the scene-aware prompt generation method introduced in Sec. 2.2.
@@ 2.2 Panoramic Data Curation Pipeline
Scene-Aware Prompt for Image Conditions. As stataed in Sec. 2.1, for image-to-panorama generation, we utilize both image and text conditions. A straightforward approach to obtain the text prompt would be to use a vision-language model (VLM) to generate a caption for the input image. However, user-provided images often contain prominent objects (e.g., a statue), and conditioning the generation process on the input image’s caption may lead to unwanted replication of these elements in the synthesized panorama (e.g., duplicate statues).
@@ 2.3 Agentic World Layering
Instance Recognition. In the context of a playable 3D world, interaction with specific scene objects is essential, while background elements typically remain static. To enable an interactive 3D scene, we must model each interactive object individually in 3D. Our approach begins with identifying which scene objects require such modeling, which demands semantic understanding and spatial relationship reasoning. Given the diversity of our generated scenes—spanning indoor/outdoor environments, natural landscapes, and game-style settings, we leverage a VLM to harness its rich world knowledge for semantic object recognition. Following instance recognition, we categorize objects into distinct sub-layers based on their semantic and spatial relationships. For instance, in an urban scene, we target segregating nearby vehicles from distant buildings into separate layers.
Layer Decomposition. Once obtaining semantic labels, the next step is to determine the precise positions of these recognized objects. However, conventional visual grounding models cannot be directly applied to panoramas due to their inherent spatial discontinuities, where objects may be fragmented across the left and right boundaries of an equirectangular projection (ERP) panorama.
@@
Layer Completion. Following object segmentation, we decompose the panorama into background (e.g., terrain) and sky layers through an autoregressive "onion-peeling" process. This involves iteratively removing recognized objects and completing each layer by inpainting occluded regions.
To train a layer completion model, we curate a panoramic object removal dataset consisting of triples of a object mask, the original panorama containing the object, and a target panorama with the object removed. Using this dataset, we fine-tune our Panorama-DiT model to learn the conditional generation of occluded regions. Similarly, we also finetune a completion model for sky layer on a dataset of sky HDRIs.
@@ 3.2 Panorama Generation
Results. The quantitative results presented in Tab. 1 demonstrate that HunyuanWorld 1.0 consistently outperforms both baseline methods across all evaluation metrics. These findings highlight the effectiveness of our method in producing high-fidelity panoramic images while preserving strong semantic correspondence with the input. Qualitative comparisons illustrated in Fig. 6 and Fig. 7 corroborate these quantitative findings. In contrast to baseline approaches, which frequently exhibit discontinuous artifacts and geometric distortions, our method generates panoramic scenes with enhanced visual coherence and aesthetic quality.
@@
Results. The quantitative results in Tab. 2 demonstrate that HunyuanWorld 1.0 achieves superior performance across all evaluation metrics compared to the baseline methods. Qualitative results in Fig. 8 and Fig. 9 reveal that our approach exhibits exceptional fidelity to textual descriptions while maintaining high visual quality standards. Furthermore, HunyuanWorld 1.0 excels at generating panoramic scenes across diverse artistic styles and thematic contexts.
@@
Results. The quantitative results presented in Tab. 3 demonstrate that HunyuanWorld 1.0 consistently outperforms both baseline methods on the visual quality of rendered novel views and semantic alignment with the input image. Qualitative comparisons illustrated in Fig. 12 show our method generates 3D worlds with superior visual quality and geometric consistency compared to the baseline approaches while maintaining enhanced alignment with the input images.
@@
Results. We present the quantitative results in Tab. 4, with qualitative comparisons shown in Fig. 13. HunyuanWorld 1.0 consistently outperforms both baseline methods across all evaluation metrics. The visual results demonstrate that our approach generates 3D worlds with high visual fidelity and strong semantic alignment with the input text descriptions. Notably, Director3D exhibits limitations in generating long-range camera trajectories for many test cases, which restricts its ability to generalize across diverse input conditions.
@@ 4 Related Work
3D World Generation. Compared to videos, 3D scene assets offer superior compatibility with standard computer graphics pipelines and ensure stronger consistency. Existing 3D world generation methods can be categorized into procedural-based methods [33, 39, 72] and learning-based methods [64, 4, 69, 68, 10, 29, 43]. Procedural generation techniques automate the creation of 3D scenes by leveraging predefined rules or constraints. Among these, rule-based methods [33, 38] employ explicit algorithms to directly generate scene geometry, laying the foundation for visual rendering. In contrast, optimization-based generation [5, 39] frames scene synthesis as an optimization problem, using cost functions based on predefined constraints (e.g., physics or design principles). A more recent paradigm, LLM-based generation (e.g., LayoutGPT [9], SceneX [72]) enables users to define environments through natural language descriptions, offering greater flexibility and user control over scene design. Learning-based approaches focus on reconstructing 3D scenes from visual inputs. These methods typically fine-tune existing diffusion models using 3D-consistent data, enabling the generation of dense multi-view images from sparse known viewpoints. They then employ per-scene 3D/4D Gaussian Splatting optimization to reconstruct the full scene. Dimension-X [43] uses video diffusion models to train separate LoRA models, decoupling spatial and temporal dimensions for multi-view video generation. LucidDreamer [4, 69, 68] generates multi-view images via progressive inpainting and multi-view diffusion models, respectively. Some methods [41, 64, 20] approach 3D scene reconstruction by first inpainting occluded regions of a panorama, then lifting 2D views to 3D Gaussian Splatting (3DGS) through scene-specific optimization. In contrast, our work targets generating layered mesh assets that can be directly integrated into existing computer graphics pipelines.
@@ 5 Conclusion
  • Core Contributors: Zhenwei Wang, Yuhao Liu, Junta Wu, Zixiao Gu, Haoyuan Wang, Xuhui Zuo, Tianyu Huang, Wenhuan Li

### v2

@@ 2.1 Generating Panoramas As World Proxy
Panorama Generation. The architecture of our panorama generation model (Panorama-DiT) is based on the Diffusion Transformer (DiT) framework [35]. For text-to-panorama generation, only the enhanced text prompts are fed into the diffusion model as condition. For image-to-panorama generation, we first project the input image into panoramic space and then encode it into the latent space using a variational autoencoder (VAE). Next, we concatenate the condition image with the noisy latent for diffusion model. To enhance the generation quality and provide additional control, the image-to-panorama generation process is also conditioned on an auxiliary textual description produced by the scene-aware prompt generation method introduced in Sec.˜2.2.
@@ 2.2 Panoramic Data Curation Pipeline
Scene-Aware Prompt for Image Conditions. As stataed in Sec.˜2.1, for image-to-panorama generation, we utilize both image and text conditions. A straightforward approach to obtain the text prompt would be to use a vision-language model (VLM) to generate a caption for the input image. However, user-provided images often contain prominent objects (e.g., a statue), and conditioning the generation process on the input image’s caption may lead to unwanted replication of these elements in the synthesized panorama (e.g., duplicate statues).
@@ 2.3 Agentic World Layering
Instance Recognition. In the context of a playable 3D world, interaction with specific scene objects is essential, while background elements typically remain static. To enable an interactive 3D scene, we must model each interactive object individually in 3D. Our approach begins with identifying which scene objects require such modeling, which demands semantic understanding and spatial relationship reasoning. Given the diversity of our generated scenes—spanning indoor/outdoor environments, natural landscapes, and game-style settings, we leverage a VLM to harness its rich world knowledge for semantic object recognition. Following instance recognition, we further categorize objects into multiple sub-layers based on their semantic and spatial relationships. For instance, in an urban scene, we target segregating nearby vehicles from distant buildings into separate layers.
Layer Decomposition. Once obtaining semantic labels, the next step is to determine the precise positions of these recognized objects. However, conventional visual grounding models cannot be directly applied to panoramas due to their inherent spatial discontinuities, where objects may be fragmented across the left and right boundaries of an equirectangular projection panorama.
@@
Layer Completion. Following object segmentation, we decompose the panorama into background (e.g., terrain) and sky layers through an autoregressive process. This involves iteratively removing recognized objects and completing each layer by inpainting occluded regions.
To train a layer completion model for panoramic scenes, we curate a panoramic object removal dataset consisting of triples of a object mask, the original panorama containing the object, and a target panorama with the object removed. Using this dataset, we fine-tune our Panorama-DiT model to learn faithful content completion of occluded regions. Similarly, we also finetune a specific completion model for sky layer on a dataset of sky HDRIs.
@@ 3.2 Panorama Generation
Results. The quantitative results presented in Tab.˜1 demonstrate that HunyuanWorld 1.0 consistently outperforms both baseline methods across all evaluation metrics. These findings highlight the effectiveness of our method in producing high-fidelity panoramic images while preserving strong semantic correspondence with the input. Qualitative comparisons illustrated in Fig.˜6 and Fig.˜7 corroborate these quantitative findings. In contrast to baseline approaches, which frequently exhibit discontinuous artifacts and geometric distortions, our method generates panoramic scenes with enhanced visual coherence and aesthetic quality.
@@
Results. The quantitative results in Tab.˜2 demonstrate that HunyuanWorld 1.0 achieves superior performance across all evaluation metrics compared to the baseline methods. Qualitative results in Fig.˜8 and Fig.˜9 reveal that our approach exhibits exceptional fidelity to textual descriptions while maintaining high visual quality standards. Furthermore, HunyuanWorld 1.0 excels at generating panoramic scenes across diverse artistic styles and thematic contexts.
@@
Results. The quantitative results presented in Tab.˜3 demonstrate that HunyuanWorld 1.0 consistently outperforms both baseline methods on the visual quality of rendered novel views and semantic alignment with the input image. Qualitative comparisons illustrated in Fig.˜12 show our method generates 3D worlds with superior visual quality and geometric consistency compared to the baseline approaches while maintaining enhanced alignment with the input images.
@@
Results. We present the quantitative results in Tab.˜4, with qualitative comparisons shown in Fig.˜13. HunyuanWorld 1.0 consistently outperforms both baseline methods across all evaluation metrics. The visual results demonstrate that our approach generates 3D worlds with high visual fidelity and strong semantic alignment with the input text descriptions. Notably, Director3D exhibits limitations in generating long-range camera trajectories for many test cases, which restricts its ability to generalize across diverse input conditions.
@@ 4 Related Work
3D World Generation. Compared to videos, 3D scene assets offer superior compatibility with standard computer graphics pipelines and ensure stronger consistency. Existing 3D world generation methods can be categorized into procedural-based methods [33, 39, 72] and learning-based methods [64, 4, 69, 68, 10, 29, 43]. Procedural generation techniques automate the creation of 3D scenes by leveraging predefined rules or constraints. Among these, rule-based methods [33, 38] employ explicit algorithms to directly generate scene geometry, laying the foundation for visual rendering. In contrast, optimization-based generation [5, 39] frames scene synthesis as an optimization problem, using cost functions based on predefined constraints (e.g., physics or design principles). A more recent paradigm, LLM-based generation (e.g., LayoutGPT [9], SceneX [72]) enables users to define environments through natural language descriptions, offering greater flexibility and user control over scene design. Learning-based approaches focus on reconstructing 3D scenes from visual inputs. These methods typically fine-tune existing diffusion models using 3D-consistent data, enabling the generation of dense multi-view images from sparse known viewpoints. They then employ per-scene 3D/4D Gaussian Splatting optimization to reconstruct the full scene. Dimension-X [43] uses video diffusion models to train separate LoRA models, decoupling spatial and temporal dimensions for multi-view video generation. Some works [4, 69, 68] generate novel-view images via progressive inpainting and world updating with depth as guidance. Some methods [41, 64, 20] approach 3D scene reconstruction by first inpainting occluded regions of a panorama, then lifting 2D views to 3D Gaussian Splatting (3DGS) through scene-specific optimization. Specifically, Layerpano3D [64] proposed to disentangle different components in a panorama with depth clustering for 3DGS reconstruction, highlighting the effectiveness of layered representation. In contrast, our work targets generating layered mesh assets that can be directly plugged into existing computer graphics workflows.
@@ 5 Conclusion
  • Core Contributors: Tengfei Wang, Zhenwei Wang, Yuhao Liu, Junta Wu, Zixiao Gu, Haoyuan Wang, Xuhui Zuo, Tianyu Huang, Wenhuan Li